OpenStudy (anonymous):

I'm torn between -3 and 3, please help. The graph of f(x) = (0.5)^x is replaced by the graph of g(x) = (0.5)^x - k. If g(x) is obtained by shifting f(x) down by 3 units, the value of k is ________.

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

I'm thinking it is just 3

OpenStudy (campbell_st):

use this site to check your choices... https://www.desmos.com/calculator it allows you to better understand the shifts graph the curve f(x) and then your g(x) and see what happens

OpenStudy (anonymous):

Doesn't really help, imo

OpenStudy (campbell_st):

ok...

OpenStudy (anonymous):

Sorry!

OpenStudy (anonymous):

But I don't think that really helps in this type of problem

OpenStudy (anonymous):

Nevermind, It's 3, I guess the graph kinda did help :D

OpenStudy (anonymous):

Because if it were -3, the graph would go upward.

OpenStudy (campbell_st):

well type in y = (0.5)^x then y = (0.5)^x - 3 you'll see 2 graphs... different colours... and you should be able to notice a difference. I just did it and it worked for me

OpenStudy (anonymous):

But with positive 3 it goes down

OpenStudy (anonymous):

Thanks :>

OpenStudy (campbell_st):

if you want + 3 do it as y = (0.5)^(x + 3)